OpenStudy (anonymous):

hi guys what about Q1 R=9 and rm=9 plz

OpenStudy (anonymous):

Sorry, buy I got wrong #1, but I will like to know wich is the answer of Q2 b)

OpenStudy (benzo):

q1 a)ia=2*i1*R/(2*R+rm) and vo=ia*rm

OpenStudy (anonymous):

q1 1)6 2)-3 3)3 4)18